Could someone please help me with this problem? I have to correct it on my homework and keep coming up with 70 which the teacher marked wrong.
-4^2 + 3^3[-2(3-2^2)]

Answer by Alan3354(69443)

-4^2 + 3^3[-2(3-2^2)]
= -16 + 27*(-2*(3-4))
= -16 + 27*(-2*-1)
= -16 + 27*(2)
= -16 + 54
= 38
-----------
If the 1st term is -4^2 = -16
If it's (-4)^2 = +16, which would give an answer of 70